# Conformal e-prediction

### v4

@@ Abstract
This paper discusses a counterpart of conformal prediction for e-values, conformal e-prediction. Conformal e-prediction is conceptually simpler and had been developed in the 1990s as precursor of conformal prediction. When conformal prediction emerged as result of replacing e-values by p-values, it seemed to have important advantages over conformal e-prediction without obvious disadvantages. This paper re-examines relations between conformal prediction and conformal e-prediction systematically from a modern perspective. Conformal e-prediction has advantages of its own, such as the ease of designing conditional conformal e-predictors and the guaranteed validity of cross-conformal e-predictors (whereas for cross-conformal predictors validity is only an empirical fact and can be broken with excessive randomization). Even where conformal prediction has clear advantages, conformal e-prediction can often emulate those advantages, more or less successfully.
@@
Soon after the publication of [5, 24], conformal e-prediction seems to have disappeared. Perhaps the main reason why it was superseded by conformal prediction was that conformal predictions can be packaged as prediction sets [25, Sect. 2.2], and in this case their property of validity is very easy to state: we just say that the probability of error is at most $\epsilon$ at a prespecified significance level $\epsilon$ [25, Proposition 2.3]. This was clearly stated only in 2001 [12, Theorem 1], although this statement was implicit in the standard requirement of validity for p-values stated in [24]. The standard requirement of validity for e-values, also stated in [24], does not admit such an intuitive restatement in terms of probability of error (see, however, Appendix B).
Another reason for conformal e-prediction losing its popularity may have been the finding in 2002 [21, Theorem 1] that, in the on-line mode of prediction, smoothed conformal predictors make errors independently. An important corollary of this stronger property of validity is that small probabilities of errors manifest themselves, with high probability, as a low frequency of errors [25, Corollary 2.5].
@@
For simplicity, in this paper we consider IID (or at least exchangeable) data and concentrate on the problem of pattern recognition (also known as classification). We start in Sect. 1 from the definition of conformal e-prediction and continue in Sect. 2 with discussing its validity. The simplest property of validity (Proposition 1 in Sect. 2) consists in conformal e-predictors producing at each step a valid e-value for the true label: namely, its expectation is at most 1 at each step; the general notion of e-value is introduced right after Proposition 1. Since the expectation is the average over the sample space, we can say that conformal e-predictors are valid in the space domain, or space-wise. A complementary notion of efficiency is efficiency in the time domain, which we consider next.
Proposition 1 does not say anything about the relation between the e-values for the true labels produced at different steps. Is it possible that for some streams of data the average of the e-values produced at different steps tends to 2 while for others it tends to 0.5 with non-zero probability? In Sect. 2 we show that this is impossible in the online prediction protocol, stating both the strong law of large numbers and the law of the iterated logarithm for the e-values produced for the true labels at different steps. This does not fully replace the strong property of independence of errors for conformal prediction, but it can be regarded as a partial replacement. We can see that the e-values produced at different steps are not misleading in the time domain; not only is their expectation at most 1 at each step, their average is at most 1 time-wise in the long run.
The properties of validity established in Sect. 2 are marginal, in the sense that the expectations, or time averages, in them are not conditional on any properties of the observations. For example, if our predictions are for people, in principle we can get very different averages for men and women. In the case of conformal prediction, a simple way of achieving conditional validity is using Mondrian conformal prediction [25, Sect. 4.6]. Mondrian conformal prediction requires a hard partition of observations, such as the partition of people into men and women. Interestingly, conformal e-prediction is much more flexible when trying to achieve conditional validity. Its conditional version does not have to be based on a partition (Sect. 3); e.g., we may require separate validity for men, women, and Europeans. It has been shown recently that this type of conditionality can also be achieved for conformal prediction (see, e.g., [6]), but it is much easier to achieve and appears to be more natural in the case of conformal e-prediction. This can be regarded as an advantage of conformal e-prediction.
@@
To prevent loss in computational efficiency without sacrificing predictive efficiency, cross-conformal predictors were introduced in [23]. Cross-conformal predictors are not provably valid [23, Appendix], and this sometimes even shows in experimental results [11]. The limits of violations of validity are given by Rüschendorf’s result (see, e.g., [27]): when merging p-values coming from different folds by taking arithmetic mean (this is essentially what cross-conformal predictors do), the resulting arithmetic mean has to be multiplied by 2 in order to guarantee validity. In the more recent method of jackknife+, introduced in [2] and closely related to cross-conformal prediction, there is a similar factor of 2 [2, Theorem 1], which cannot be removed in general [2, Theorem 2].
In Sect. 5, we introduce a version of cross-conformal prediction based on e-values, which we call cross-conformal e-prediction. The situation with cross-conformal e/prediction is very different from cross-conformal prediction, as the arithmetic mean of e-values is always an e-value. This is an obvious fact, and it is shown in [28, Sect. 3] that arithmetic mean is the only useful merging rule. Therefore, cross-conformal e/prediction is always valid. This is a second advantage of conformal e-prediction.
The emphasis of Sections 25 is on the validity of conformal e-prediction and its computational efficiency, while Sect. 6 moves on to its predictive efficiency. What are suitable criteria of predictive efficiency? We propose two such criteria in the case of pattern recognition, the “observed log criterion” and the “prior log criterion”.
@@ 1 Conformal e-predictors
Suppose we are given a training set $z_{1},\dots,z_{n}$ consisting of labelled objects $z_{i}=(x_{i},y_{i})$ and our goal is to predict the label of a new object $x$. In this paper we consider predictors of the following type: for each potential label $y$ for $x$ we would like to have a number $f(z_{1},\dots,z_{n},x,y)$ reflecting the plausibility of $y$ being the true label of $x$. An example is conformal transducers [25, Sect. 2.7], which, in the terminology of this paper, may be called conformal p-predictors. The output
@@
A nonconformity e-measure is a measurable function $A:\mathbf{Z}^{+}\to[0,\infty)^{+}$ that maps any finite sequence $(z_{1},\dots,z_{m})$, $m\in\{1,2,\dots\}$, to a finite sequence $(\alpha_{1},\dots,\alpha_{m})$ of the same length consisting of nonnegative numbers with average at most 1,
@@
The conformal e-predictor $f$ corresponding to such $A$ is defined by
@@
$\left(f(z_{1},\dots,z_{n},x,y)\mid y\in\mathbf{Y}\right).$
The full prediction for the label of $x$ can be summarized as, e.g., the point prediction
@@
and the e-credibility $f(z_{1},\dots,z_{n},x,\hat{y})$. We can make a confident point prediction when the e-confidence is large while the e-credibility is not.
Let us say that a nonconformity e-measure and the corresponding conformal e-predictor are admissible if we always have “$=$” in place of $\leq$” in the definition (1). If a conformal e-predictor is not admissible, we can make its predictions more confident without sacrificing their validity. Therefore, we will usually concentrate on admissible nonconformity e-measure and admissible conformal e-predictors.
A nonnegative nonconformity measure $A:\mathbf{Z}^{+}\to[0,\infty)^{+}$ is defined as a nonconformity e-measure except that the condition (1) is omitted. Given a nonnegative nonconformity measure $A$, we can always define the corresponding nonconformity e-measure $A^{\prime}$ by normalizing $A$:
@@
A further generalization of nonnegative nonconformity measures, nonconformity measures, is obtained by dropping the condition of nonnegativity; these are functions $A:\mathbf{Z}^{+}\to\mathbb{R}$. They are used, explicitly or implicitly, in conformal p-prediction.
@@
$f(z_{1},\dots,z_{n+1}):=\begin{cases}(n+1)/\left|\textrm{SV}\right|&\text{if $n+1\in\textrm{SV}$}\\ 0&\text{otherwise},\end{cases}$
where SV is the set of indices of support vectors: $i\in\textrm{SV}$ if and only if $z_{i}$, $i\in\{1,\dots,n+1\}$, is a support vector for the SVM constructed from $z_{1},\dots,z_{n+1}$ as training set. It is based on the indicator function of being a support vector. When given a training set $z_{1},\dots,z_{n}$ and a new object $x$, this conformal e-predictor goes through all potential labels $y$ for $x$ and for each constructs an SVM and outputs $f(z_{1},\dots,z_{n},x,y)$. It makes it computationally inefficient.
@@ 2 Validity of conformal e-predictors
The following obvious proposition asserts the validity of conformal e-predictors. Let us write $Z_{1},Z_{2},\dots$ for the random vectors whose realizations are the observed data $z_{1},z_{2},\dots$; more generally, $(X,Y)$ or $Z$ are random vectors with values in the observation space $\mathbf{Z}$. As usual, a finite sequence of random vectors is exchangeable if its joint distribution does not change if it is permuted (and an infinite sequence is exchangeable if its joint distribution does not change if its first $n$ elements are permuted, for any $n$). The difference between exchangeability and being IID (independent and identically distributed) can be substantial for finite data sequences (but for infinite data sequences the difference between the two assumptions disappears according to de Finetti’s theorem [25, Sect. A.5.1]).
@@
The property of validity given in Proposition 1 says that conformal e-predictors output valid e-values for the true labels. Formally, an e-variable is a random variable $E$ satisfying $\mathbb{E}(E)\leq 1$ under the true data-generating distribution (and the values it takes are e-values). We do not expect the e-values for the true labels to be large because, by Markov’s inequality, $\mathbb{P}(E\geq C)\leq 1/C$ for any constant $C>1$.
Proposition 1 involves a “space average”, i.e., an average over the sample space. The analogous property of validity for conformal prediction can be stated in terms of error probabilities for set predictions (see, e.g., [25, Proposition 2.1]), which is particularly intuitive. This is the first advantage of conformal prediction pointed out in Sect. Conformal e-prediction. It disappears when we move to conformal e-prediction: validity has to be defined in terms of e-values (however, it has been argued [16] that e-values are more intuitive than p-values).
Another advantage of conformal prediction is that, in the online mode of prediction (to be defined shortly), conformal predictors (in their smoothed version) make errors at different steps independently (see, e.g., [25, Theorem 11.1]). Without independence, it is possible, even when the probability of error at each step is $\epsilon$, for the long-term relative frequency of errors over consecutive steps to be either 0 or 1 (1 with probability $\epsilon$). The independence of errors forces the long-term frequency of errors to be $\epsilon$ almost surely (which is also asserted in [25, Proposition 2.1]). We can say that independence implies ergodicity, i.e., the almost sure coincidence of time and space averages.
@@
To give an example demonstrating the loss of independence in conformal e-prediction, we first define the online prediction protocol. In the online protocol for conformal e-prediction, we observe an object $x_{1}$, apply the conformal e/predictor to compute the conformal e-values $f(x_{1},y)$ for all possible labels $y\in\mathbf{Y}$, observe the true label $y_{1}$, record the e-value $e_{1}:=f(x_{1},y_{1})$ for it, observe another object $x_{2}$, apply the conformal e-predictor to compute the e-values $f(x_{1},y_{1},x_{2},y)$ for all possible labels $y\in\mathbf{Y}$, observe the true label $y_{2}$, record the e-value $e_{2}:=f(x_{1},y_{1},x_{2},y_{2})$ for it, etc. In the case conformal prediction, we can instead record conformal p-values $p_{n}$ (or record whether a mistake is made at a given significance level). While the conformal p-values are independent (for smoothed conformal predictors), the following example shows that the conformal e-values are not independent already in a toy binary situation (perhaps this is the simplest non-trivial example of conformal e-prediction). We write $E_{n}$ for the e-value $e_{n}$ regarded as a random variable (an e-variable).
@@
We follow the proof of [22, Lemma 14] (given as Lemma 3.15 in the first edition of [25]) and use the terminology of [19, Chap. 7].
Let $\mathcal{F}_{n}$ be the $\sigma$-algebra generated by the multiset $\lbag Z_{1},\dots,Z_{n-1}\rbag$ and the observations $Z_{n},Z_{n+1},\dots$. (These $\sigma$-algebras form the exchangeable filtration [15, Sect. 5.6].) For each time horizon $N\in\{2,3,\dots\}$, the stochastic sequence $(E_{n}-1,\mathcal{F}_{n})$, $n=N,\dots,1$, is a bounded supermartingale difference. By Hoeffding’s inequality (see, e.g., [25, Sect. A.6.3]), for any $N\in\{2,3,\dots\}$,
@@
where $C$ is an upper bound on the conformal e-predictor $f$. By the Borel–Cantelli lemma [18, Sect. 2.10, part (a) of the lemma], the internal inequality in (7) holds only for finitely many $N$ for a fixed $\epsilon>0$. This implies (6).
@@
The last inequality says that $E_{n}\wedge N^{1/3}$ are approximate time-wise e-values (assuming that $\epsilon$ is small and $N\gg\epsilon^{-6}$), and so $E_{n}$ are approximate time-wise e-values if we regard an e-value of $N^{1/3}$ as large enough for the difference between $N^{1/3}$ and larger e-values to be considered unimportant (such as an e-value of 100 on Jeffreys’s scale [9, Appendix B]). (To derive (9) from (8), just set $C:=N^{1/3}$.)
@@
(where $\delta_{a}$ is the probability measure concentrated on $\{a\}$), since $E_{N},\dots,E_{1}$ are generated independently. This gives us a product distribution for $E_{1},E_{2},\dots$, namely (10) with $N$ replaced by $\infty$. For any $b\in\{1,2,\dots\}$, by Le Cam’s version of Poisson’s theorem [18, Sect. 3.12], the probability of the event that $b$ of the $E_{n}$ for $n\in[2^{k},2^{k+1})$ will satisfy $E_{n}=n$ tends to $e^{-\lambda}\lambda^{b}/b!$ as $k\to\infty$, where $\lambda:=\ln 2$; therefore, this event will happen infinitely often (for any $b$). Therefore, (6) will be violated almost surely; namely, the limsup in it will be $\infty$ almost surely. Figure 1 illustrates this by plotting the average conformal e-value $\frac{1}{N}\sum_{n=1}^{N}e_{n}$ vs $N$ for a wide range of $N$, $N\in\{1,\dots,10^{10}\}$ (with $N$ given on a logarithmic scale).
@@
In conclusion of this section, let me state what I consider to be the main property of validity for e-conformal prediction (it implies the other properties listed here, although by itself it is not particularly intuitive): $\mathbb{E}(E_{n}\mid\mathcal{F}_{n+1})\leq 1$, where $\mathcal{F}$ is the exchangeable filtration and $E_{n}$ is a conformal e/variable. It was introduced in the proof of Proposition 3 but is worth stating separately.

### v5

# wideformula*

@@ Abstract
This paper discusses a counterpart of conformal prediction for e-values, conformal e-prediction. Conformal e-prediction is conceptually simpler and had been developed in the 1990s as a precursor of conformal prediction. When conformal prediction emerged as result of replacing e-values by p-values, it seemed to have important advantages over conformal e-prediction without obvious disadvantages. This paper re-examines relations between conformal prediction and conformal e-prediction systematically from a modern perspective. Conformal e-prediction has advantages of its own, such as the ease of designing conditional conformal e-predictors and the guaranteed validity of cross-conformal e-predictors (whereas for cross-conformal predictors validity is only an empirical fact and can be broken with excessive randomization). Even where conformal prediction has clear advantages, conformal e-prediction can often emulate those advantages, more or less successfully.
@@ 1 Introduction
Soon after the publication of [6, 25], conformal e-prediction seems to have disappeared. Perhaps the main reason why it was superseded by conformal prediction was that conformal predictions can be packaged as prediction sets [26, Sect. 2.2], and in this case their property of validity is very easy to state: we just say that the probability of error is at most $\epsilon$ at a prespecified significance level $\epsilon$ [26, Proposition 2.3]. This was clearly stated only in 2001 [13, Theorem 1], although this statement was implicit in the standard requirement of validity for p-values stated in [25]. The standard requirement of validity for e-values, also stated in [25], does not admit such a simple restatement in terms of probability of error without weakening it drastically; see Appendix B for further details. (While the prediction sets derived from conformal e-prediction can be used to define the property of validity in its strong form, validity becomes a property of the whole family of prediction sets for different significance levels.)
Another reason for conformal e-prediction losing its popularity may have been the finding in 2002 [22, Theorem 1] that, in the on-line mode of prediction, smoothed conformal predictors make errors independently. An important corollary of this stronger property of validity is that small probabilities of errors manifest themselves, with high probability, as a low frequency of errors [26, Corollary 2.5].
@@
For simplicity, in this paper we consider IID (or at least exchangeable) data and concentrate on the problem of pattern recognition (also known as classification). We start in Sect. 2 from the definition of conformal e-prediction and continue in Sect. 3 with discussing its validity. The simplest property of validity (Proposition 2 in Sect. 3) consists in conformal e-predictors producing at each step a valid e-value for the true label: namely, its expectation is at most 1 at each step; the general notion of e-value is introduced right after Proposition 2. Since the expectation is the average over the sample space, we can say that conformal e-predictors are valid in the space domain, or space-wise. A complementary notion of efficiency is efficiency in the time domain, which we consider next.
Proposition 2 does not say anything about the relation between the e-values for the true labels produced at different steps. Is it possible that for some streams of data the average of the e-values produced at different steps tends to 2 while for others it tends to 0.5 with non-zero probability? In Sect. 3 we show that this is impossible in the online prediction protocol, stating both the strong law of large numbers and the law of the iterated logarithm for the e-values produced for the true labels at different steps. This does not fully replace the strong property of independence of errors for conformal prediction, but it can be regarded as a partial replacement. We can see that the e-values produced at different steps are not misleading in the time domain; not only is their expectation at most 1 at each step, their average is at most 1 time-wise in the long run.
The properties of validity established in Sect. 3 are marginal, in the sense that the expectations, or time averages, in them are not conditional on any properties of the observations. For example, if our predictions are for people, in principle we can get very different averages for men and women. In the case of conformal prediction, a simple way of achieving conditional validity is using Mondrian conformal prediction [26, Sect. 4.6]. Mondrian conformal prediction requires a hard partition of observations, such as the partition of people into men and women. Interestingly, conformal e-prediction is much more flexible when trying to achieve conditional validity. Its conditional version does not have to be based on a partition (Sect. 4); e.g., we may require separate validity for men, women, and Europeans. It has been shown recently that this type of conditionality can also be achieved for conformal prediction (see, e.g., [7]), but it is much easier to achieve and appears to be more natural in the case of conformal e-prediction. This can be regarded as an advantage of conformal e-prediction.
@@
To prevent loss in computational efficiency without sacrificing predictive efficiency, cross-conformal predictors were introduced in [24]. Cross-conformal predictors are not provably valid [24, Appendix], and this sometimes even shows in experimental results [12]. The limits of violations of validity are given by Rüschendorf’s result (see, e.g., [28, Proposition 2]): when merging p-values coming from different folds by taking arithmetic mean (this is essentially what cross-conformal predictors do), the resulting arithmetic mean has to be multiplied by 2 in order to guarantee validity. In the more recent method of jackknife+, introduced in [3] and closely related to cross-conformal prediction, there is a similar factor of 2 [3, Theorem 1], which cannot be removed in general [3, Theorem 2].
In Sect. 6, we introduce a version of cross-conformal prediction based on e-values, which we call cross-conformal e-prediction. The situation with cross-conformal e/prediction is very different from cross-conformal prediction, as the arithmetic mean of e-values is always an e-value. This is an obvious fact, and it is shown in [29, Sect. 3] that arithmetic mean is the only useful merging rule. Therefore, cross-conformal e/prediction is always valid. This is a second advantage of conformal e-prediction.
The emphasis of Sections 36 is on the validity of conformal e-prediction and its computational efficiency, while Sect. 7 moves on to its predictive efficiency. What are suitable criteria of predictive efficiency? We propose two such criteria in the case of pattern recognition, the “observed log criterion” and the “prior log criterion”.
@@ 2 Conformal e-predictors
Suppose we are given a training set $z_{1},\dots,z_{n}$ consisting of labelled objects $z_{i}=(x_{i},y_{i})$ and our goal is to predict the label of a new object $x$. In this paper we consider predictors of the following type: for each potential label $y$ for $x$ we would like to have a number $f(z_{1},\dots,z_{n},x,y)$ reflecting the plausibility of $y$ being the true label of $x$. An example is conformal transducers [26, Sect. 2.7], which, in the terminology of this paper, may be called conformal p-predictors. The output
@@
A nonconformity e-measure is a measurable function $A:\mathbf{Z}^{+}\to[0,\infty)^{+}$ that maps any finite sequence $(z_{1},\dots,z_{m})$, $m\in\{1,2,\dots\}$, to a finite sequence $(\alpha_{1},\dots,\alpha_{m})$ of the same length consisting of nonnegative numbers (nonconformity scores) with average at most 1,
@@
Very roughly, this property means that each nonconformity score $\alpha_{i}$ (supposed to measure the strangeness of $z_{i}$ as compared with the other observations in the sequence) does not depend on the position of $z_{i}$, or of the other observations, in the sequence. The conformal e-predictor $f$ corresponding to such $A$ is defined by
@@
$\left(f(z_{1},\dots,z_{n},x,y)\mid y\in\mathbf{Y}\right).$ (4)
We can regard the family (4) of e-values for each potential label $y$ as a soft set predictor. By thresholding $f$ at some level, we can get a (hard) set predictor (as in (42) below); we include in the prediction set for the label of $x$ the potential labels $y$ for which the conformal e-value is less than a chosen level. But there is no need to choose the level in advance, and we interpret the conformal e-value $f(z_{1},\dots,z_{n},x,y)$ of $y$ as the degree to which $y$ is excluded from the soft prediction set. We want our conformal prediction to be valid and efficient, where validity means that we do not want the true label to be excluded (i.e., to have a large e-value) while efficiency means that we want to exclude all other labels.
The full prediction (4) for the label of $x$ can be summarized as, e.g., the point prediction
@@
and the e-credibility $f(z_{1},\dots,z_{n},x,\hat{y})$. (See [26, Sect. 3.5.1] for their p-counterparts.) We can make a confident point prediction when the e-confidence is large while the e-credibility is not.
Let us say that a nonconformity e-measure and the corresponding conformal e-predictor are admissible if we always have “$=$” in place of $\leq$” in the definition (2). If a conformal e-predictor is not admissible, we can make its predictions more confident without sacrificing their validity. Therefore, we will usually concentrate on admissible nonconformity e-measures and admissible conformal e-predictors.
A nonnegative nonconformity measure $A:\mathbf{Z}^{+}\to[0,\infty)^{+}$ is defined as a nonconformity e-measure except that the condition (2) is omitted. Given a nonnegative nonconformity measure $A$, we can always define the corresponding admissible nonconformity e-measure $A^{\prime}$ by normalizing $A$:
@@
A further generalization of nonnegative nonconformity measures, nonconformity measures, is obtained by dropping the condition of nonnegativity; these are functions $A:\mathbf{Z}^{+}\to\mathbb{R}^{+}$. They are used, explicitly or implicitly, in conformal p-prediction. See, e.g., [1, Sect. 1.3] (and Remark 1 below).
@@
$f(z_{1},\dots,z_{n+1}):=\begin{cases}(n+1)/\left|\textrm{SV}\right|&\text{if $n+1\in\textrm{SV}$}\\ 0&\text{otherwise},\end{cases}$ (6)
where SV is the set of indices of support vectors: $i\in\textrm{SV}$ if and only if $z_{i}$, $i\in\{1,\dots,n+1\}$, is a support vector for the SVM constructed from $z_{1},\dots,z_{n+1}$ as training set. It is based on the indicator function of being a support vector; indeed, the right-hand side of (5) then becomes
$\left(\frac{(n+1)1_{\{i\in\textrm{SV}\}}}{\sum_{i=1}^{n+1}1_{\{i\in\textrm{SV}\}}}\right)_{i=1}^{n+1},$
which agrees with (6). When given a training set $z_{1},\dots,z_{n}$ and a new object $x$, this conformal e-predictor goes through all potential labels $y$ for $x$ and for each constructs an SVM and outputs $f(z_{1},\dots,z_{n},x,y)$. It makes it computationally inefficient.
Remark 1.
In conformal prediction, several natural ways of defining nonconformity measures have been discussed and widely used in literature. The definition at the beginning of this section parallels the definition of nonconformity measures in [1, Sect. 1.3]; under these definitions, nonconformity measures map sequences of observations to sequences of nonconformity scores. Under other definitions, nonconformity measures may map, e.g., a bag of observations and another observation $z$ to one nonconformity score (that for $z$), and there are several varieties of such definitions; see, e.g., [26, Sect. 2.9.3]. Each of these varieties could have been adapted to conformal e-prediction.
@@ 3 Validity of conformal e-predictors
The following obvious proposition asserts the validity of conformal e-predictors. Let us write $Z_{1},Z_{2},\dots$ for the random elements whose realizations are the observed data $z_{1},z_{2},\dots$; more generally, $(X,Y)$ or $Z$ are random elements with values in the observation space $\mathbf{Z}$. As usual, a finite sequence of random elements is exchangeable if its joint distribution does not change if it is permuted (and an infinite sequence is exchangeable if its joint distribution does not change if its first $n$ elements are permuted, for any $n$). The difference between exchangeability and being IID (independent and identically distributed) can be substantial for finite data sequences (but for infinite data sequences the difference between the two assumptions disappears, provided $\mathbf{Z}$ is a standard Borel space, according to de Finetti’s theorem [26, Sect. A.5.1]).
@@
The property of validity given in Proposition 2 says that conformal e-predictors output valid e-values for the true labels. Formally, an e-variable is a random variable $E$ satisfying $\mathbb{E}(E)\leq 1$ under the data-generating distribution (and the values it takes are e-values). We do not expect the e-values for the true labels to be large because, by Markov’s inequality, $\mathbb{P}(E\geq C)\leq 1/C$ for any constant $C>1$.
Proposition 2 involves a “space average”, i.e., an average over the sample space. The analogous property of validity for conformal prediction can be stated in terms of error probabilities for set predictions (see, e.g., [26, Proposition 2.1]), which is particularly intuitive. This is the first advantage of conformal prediction pointed out in Sect. 1. It disappears when we move to conformal e-prediction: validity has to be defined in a more complicated way in order to avoid making this property much weaker (however, it has been argued [17] that e-values are more intuitive than p-values).
Another advantage of conformal prediction is that, in the online mode of prediction (to be defined shortly), conformal predictors (in their smoothed version) make errors at different steps independently (see, e.g., [26, Theorem 11.1]). Without independence, it is possible, even when the probability of error at each step is $\epsilon$, for the long-term relative frequency of errors over consecutive steps to be either 0 or 1 (1 with probability $\epsilon$). The independence of errors forces the long-term frequency of errors to be $\epsilon$ almost surely (which is also asserted in [26, Proposition 2.1]). We can say that independence implies ergodicity, i.e., the almost sure coincidence of time and space averages.
@@
To give an example demonstrating the loss of independence in conformal e-prediction, we first define the online prediction protocol. In the online protocol for conformal e-prediction, we observe an object $x_{1}$, apply the conformal e/predictor to compute the conformal e-values $f(x_{1},y)$ for all possible labels $y\in\mathbf{Y}$, observe the true label $y_{1}$, record the e-value $e_{1}:=f(x_{1},y_{1})$ for it, observe another object $x_{2}$, apply the conformal e-predictor to compute the e-values $f(x_{1},y_{1},x_{2},y)$ for all possible labels $y\in\mathbf{Y}$, observe the true label $y_{2}$, record the e-value $e_{2}:=f(x_{1},y_{1},x_{2},y_{2})$ for it, etc. In the case of conformal prediction, we can instead record conformal p-values $p_{n}$ (or record whether a mistake is made at a given significance level). While the conformal p-values are independent (for smoothed conformal predictors), the following example shows that the conformal e-values are not independent already in a toy binary situation (perhaps this is the simplest non-trivial example of conformal e-prediction, albeit it is not particularly interesting per se). We write $E_{n}$ for the e-value $e_{n}$ regarded as a random variable (an e-variable).
@@
We follow the proof of [23, Lemma 14] (given as Lemma 3.15 in the first edition of [26]) and use the terminology of [20, Chap. 7].
Let $\mathcal{F}_{n}$ be the $\sigma$-algebra generated by the multiset $\lbag Z_{1},\dots,Z_{n-1}\rbag$ and the observations $Z_{n},Z_{n+1},\dots$. (These $\sigma$-algebras form the exchangeable filtration [16, Sect. 5.6].) For each time horizon $N\in\{2,3,\dots\}$, the stochastic sequence $(E_{n}-1,\mathcal{F}_{n})$, $n=N,\dots,1$, is a bounded supermartingale difference. By Hoeffding’s inequality (see, e.g., [26, Sect. A.6.3]), for any $N\in\{2,3,\dots\}$,
@@
where $C$ is an upper bound on the conformal e-predictor $f$. By the Borel–Cantelli lemma [19, Sect. 2.10, part (a) of the lemma], the internal inequality in (10) holds only for finitely many $N$ for a fixed $\epsilon>0$. This implies (9).
@@
The last inequality says that $E_{n}\wedge N^{1/3}$ are approximate time-wise e-values (assuming that $\epsilon$ is small and $N\gg\epsilon^{-6}$), and so $E_{n}$ are approximate time-wise e-values if we regard an e-value of $N^{1/3}$ as large enough for the difference between $N^{1/3}$ and larger e-values to be considered unimportant (such as an e-value of 100 on Jeffreys’s scale [10, Appendix B]). (To derive (12) from (11), just set $C:=N^{1/3}$.)
@@
(where $\delta_{a}$ is the probability measure concentrated on $\{a\}$), since $E_{N},\dots,E_{1}$ are generated independently. This gives us a product distribution for $E_{1},E_{2},\dots$, namely (13) with $N$ replaced by $\infty$. For any $b\in\{1,2,\dots\}$, by Le Cam’s version of Poisson’s theorem [19, Sect. 3.12], the probability of the event that $b$ of the $E_{n}$ for $n\in[2^{k},2^{k+1})$ will satisfy $E_{n}=n$ tends to $e^{-\lambda}\lambda^{b}/b!$ as $k\to\infty$, where $\lambda:=\ln 2$; therefore, this event will happen infinitely often (for any $b$). Therefore, (9) will be violated almost surely; namely, the limsup in it will be $\infty$ almost surely. Figure 1 illustrates this by plotting the average conformal e-value $\frac{1}{N}\sum_{n=1}^{N}e_{n}$ vs $N$ for a wide range of $N$, $N\in\{1,\dots,10^{10}\}$ (with $N$ given on a logarithmic scale).
@@
In conclusion of this section, let me state what can be considered to be the main property of validity for conformal e-prediction (it implies the other properties listed here, although by itself it is not particularly intuitive): $\mathbb{E}(E_{n}\mid\mathcal{F}_{n+1})\leq 1$, where $\mathcal{F}$ is the exchangeable filtration and $E_{n}$ is a conformal e/variable. It was introduced in the proof of Proposition 4 but is worth stating separately.